# Channel Estimation for XL-MIMO Systems with Decentralized Baseband Processing: Integrating Local Reconstruction with Global Refinement

### v2

@@ II-B Problem Description
where $\varrho>0$ denotes a regularization factor, $\mathbf{H}_{\mathrm{T}}$ represents the angular-delay channel, and $\mathbf{F}_{\mathrm{A}}\in\mathbb{C}^{N_{\mathrm{t}}\times N_{\mathrm{t}}}$ and $\mathbf{F}_{\mathrm{D}}\in\mathbb{C}^{K\times K}$ are the discrete Fourier transform (DFT) matrices of dimensions $N{\mathrm{t}}$ and $K$, respectively. Up to now, several algorithms have been developed to address the problem in (9), including the orthogonal matching pursuit (OMP) method [29] and Bayesian inference-based approaches [30, 31, 32]. However, with the increasing number of antennas and subcarriers in mmWave XL-MIMO systems, these centralized algorithms become computationally prohibitive [19]. For example, the sparse Bayesian learning (SBL) algorithm incurs a computational complexity of $\mathcal{O}(M^{3}Q^{3}K^{3})$ due to the need for matrix inversions, making it unsuitable for large-scale scenarios.
@@ III Proposed Two-Stage Estimation Scheme
In this section, we propose a two-stage reconstruction-then-refinement estimation scheme, as illustrated in Fig. 2, which achieves accuracy comparable to centralized methods while ensuring low computational complexity.
@@ III-A Stage I: Local Sparse Channel Reconstruction
$\mathbf{H}_{m}=\mathbf{F}_{\mathrm{L}}^{\mathrm{H}}\mathbf{H}_{\mathrm{T},m}^{\mathrm{L}}\mathbf{F}_{\mathrm{D}},$ (10)
@@
$\mathbf{y}_{m}=\boldsymbol{\Phi}_{m}\mathbf{x}_{m}+\mathbf{n}_{m},$ (11)
where $\mathbf{y}_{m}=\mathrm{vec}({\mathbf{Y}_{m}})\in\mathbb{C}^{QK\times 1}$ and $\mathbf{x}_{m}=\mathrm{vec}\left(\mathbf{H}_{\mathrm{T},m}^{\mathrm{L}}\right)\in\mathbb{C}^{NK\times 1}$ are the stacked received signal and local angular-delay channel, respectively. ${\boldsymbol{\Phi}}_{m}=\mathbf{F}_{\mathrm{D}}^{\mathrm{T}}\otimes\left(\mathbf{W}_{m}\mathbf{F}_{\mathrm{L}}^{\mathrm{H}}\right)\in\mathbb{C}^{QK\times NK}$ denotes the equivalent measurement matrix. Additionally, $\mathbf{n}_{m}=\mathrm{vec}({\mathbf{N}_{m}})\sim\mathcal{CN}(\mathbf{n}_{m};0,\zeta_{m}^{-1}\mathbf{I}_{QK\times QK})$ denotes the additional white Gaussian noise (AWGN) with covariance $\zeta_{m}^{-1}\mathbf{I}_{QK\times QK}$, where $\zeta_{m}$ is the unknown noise precision parameter to be estimated as well. Without loss of generality, a conjugated prior probability distribution is assigned for ${\zeta}_{m}$, i.e., $p(\zeta_{m})\propto\zeta_{m}^{-1}$.
@@
$\hat{\mathbf{x}}_{m}\propto\arg\max_{\mathbf{x}_{m}}\quad p(\mathbf{y}_{m}|\mathbf{x}_{m},\zeta_{m})p(\mathbf{x}_{m})p(\zeta_{m}),$ (12)
where $p(\mathbf{y}_{m}|\mathbf{x}_{m},\zeta_{m})$ and ${p(\mathbf{x}_{m})}$ denote the joint likelihood function of $\mathbf{x}_{m}$ and $\zeta_{m}$ and the prior probability distribution of $\mathbf{x}_{m}$, respectively. Notably, the estimation performance of problem (12) is dependent on the accuracy of prior distribution $p(\mathbf{x}_{m})$. Meanwhile, due to the limited computational capability of LPUs, a computationally efficient estimation scheme is needed. To solve this problem, we proposed a SBL-GNNs algorithm, which will be elaborated on in Sec. IV.
@@ III-B1 Channel Fusion in Angular Domain
The purpose of the central fusion is to align the local estimate $\hat{\mathbf{H}}_{\mathrm{T},m}^{\mathrm{L}}$ into the global angular domain with enhanced resolution, thereby facilitating to leverage the correlations of subarray channel in the angular-delay domain. Specifically, leveraging the linear property of the DFT transformation matrix, we have
$\mathbf{H}_{\mathrm{T}}=\sum_{m=1}^{M}\mathbf{F}_{\mathrm{A},m}\mathbf{H}_{m}\mathbf{F}_{\mathrm{D}}^{\mathrm{H}}\triangleq\sum_{m=1}^{M}\mathbf{H}_{\mathrm{T},m},$ (13)
where $\mathbf{H}_{\mathrm{T},m}$ denotes the angular-delay channels of subarray $m$ under the global angular transformation matrix $\mathbf{F}_{\mathrm{A}}=\left[\mathbf{F}_{\mathrm{A},1},\cdots,\mathbf{F}_{\mathrm{A},m},\cdots,\mathbf{F}_{\mathrm{A},M}\right]$. Based on this, the aggregated channel in the CPU can be given by
@@
Compared to the local angular-delay channels $\mathbf{H}_{\mathrm{T},m}^{\mathrm{L}}$, the global angular-delay channel $\mathbf{H}_{\mathrm{T}}$ can more efficiently capture the inherent correlations among subarrays. Specifically, as the subarrays are spatially close to one another, $\mathbf{H}_{\mathrm{T},m}$ shares common angular and delay sparse supports, and exhibits a high degree of correlation. By effectively exploiting these common sparse supports and correlations, the reconstruction performance can be significantly enhanced.
@@ III-B2 Bayesian Refinement
where $p(\mathbf{r}|\mathbf{x},\kappa)$ denotes the joint likelihood function of $\mathbf{x}$ and $\kappa$; $p(\mathbf{x})$ and $p(\kappa)\propto\kappa^{-1}$ denote the prior probability distribution of $\mathbf{x}$ and $\kappa$, respectively. To solve problem (16) efficiently, we propose a variational message-passing-based denoising algorithm that leverages a Markov chain-based hierarchical prior to capture the sparsity and correlation of $\mathbf{x}$, as detailed in Sec. V. Once the global angular-delay channel $\tilde{\mathbf{x}}$ is obtained, the global antenna-frequency-domain channel can be reconstructed as $\tilde{\mathbf{H}}=\mathbf{F}_{\mathrm{A}}^{\mathrm{F}}\tilde{\mathbf{H}}_{\mathrm{T}}\mathbf{F}_{\mathrm{D}}$ = $[\tilde{\mathbf{H}}_{1}^{\mathrm{T}},\tilde{\mathbf{H}}_{2}^{\mathrm{T}},\cdots,\tilde{\mathbf{H}}_{M}^{\mathrm{T}}]$. Then, $\tilde{\mathbf{H}}_{m}$ are sent to the corresponding LPUs for subsequent algorithm designs.
@@ IV-A SBL Estimator
$p(\mathbf{x}_{m}|\boldsymbol{\gamma}_{m})$ $=\prod_{n=1}^{N}\mathcal{CN}(x_{m,n};0,\gamma_{m,n}^{-1}),$ (17)
$p(\boldsymbol{\gamma}_{m})$ $=\prod_{n=1}^{N}p(\gamma_{m,n};\boldsymbol{\lambda}_{m,n}),$ (18)
where $\boldsymbol{\gamma}_{m}=[\gamma_{m,1},\gamma_{m,2},\cdots,\gamma_{m,N}]^{\mathrm{T}}$ and $\boldsymbol{\lambda}_{m,n}$ denote the precision parameter of variance for $x_{m,n}$ and prior distribution parameters of ${\gamma}_{m,n}$. Utilizing (17), (18) and Expectation-Maximization (EM) framework, problem (12) can be efficiently solved by SBL-based algorithms [31, 32], and the update steps in each iteration are as follows:
@@
$\boldsymbol{\lambda}_{m,n}^{(t)}$ $=g_{2}\left(\boldsymbol{\mu}_{m}^{(t-1)},\boldsymbol{\Sigma}_{m}^{(t-1)},\boldsymbol{\lambda}_{m,n}^{(t-1)}\right),$ (22)
@@
where $\boldsymbol{\Sigma}_{m}^{(t)}$ and $\boldsymbol{\mu}_{m}^{(t)}$ represent the posterior covariance matrix and mean of $\mathbf{x}_{m}$ at the $t$-th iteration, respectively, while $g_{1}(\cdot)$, $g_{2}(\cdot)$, and $g_{3}(\cdot)$ denote three non-linear mapping functions. Recently, various approaches have been proposed for designing $g_{1}(\cdot)$, $g_{2}(\cdot)$, and $g_{3}(\cdot)$, including mean-field-based methods [31] and variational inference techniques [32]. These modules execute iteratively until convergence, at which point the posterior mean $\boldsymbol{\mu}_{m}^{(t)}$ is adopted as the final estimate of $\mathbf{x}_{m}$.
Revisiting (19)-(23), the estimation performance of SBL-based estimator is highly dependent on the accuracy of the prior model. However, conventional sparse prior models often fail to fully leverage the correlations among channel entries. For instance, the Bernoulli-Gaussian prior [33] and two-layer nested Gaussian-Gamma prior [34] assume independence among channel entries, which leads to performance degradation when addressing block-sparse signals. Although advanced approaches such as the Markov chain-based and the 4-connected MRF-based hierarchical priors [10, 11, 31, 32] have been developed, these methods only consider correlations among adjacent entries. This limitation renders them inadequate for effectively capturing the global correlations of local angular-delay channels.
@@ IV-B Proposed SBL-GNNs Estimator
GNNs have recently emerged as powerful tools for statistical inference, offering a flexible framework for learning efficient message-passing strategies. Motivated by their potential, we propose a GNN-based approach within the SBL framework to estimate the prior variance of the angular-delay channel and associated parameters. Specifically, the non-linear mapping functions $g_{1}(\cdot)$ and $g_{2}(\cdot)$ in M-step are replaced with GNNs. To facilitate this, the prior variance update is formulated as a MAP estimation problem, modeled using a pairwise MRF. Leveraging the characteristics of the pairwise MRF, a GNN architecture is then designed to efficiently capture the dependencies among channel coefficients and perform MAP inference, thereby improving channel estimation accuracy and robustness.
According to SBL principles, the output of the observation module in the $t$-th iteration can be equivalently characterized as an additional white Gaussian noise (AWGN) model, i.e.,
$\boldsymbol{\mu}_{m}^{(t)}=\mathbf{x}_{m}^{(t)}+\mathbf{w}_{m}^{(t)},$ (24)
@@
$p(\mathbf{x}_{m},\boldsymbol{\gamma}_{m},\mathbf{s}_{m})$ $=\underbrace{p(\mathbf{x}_{m}|\boldsymbol{\gamma}_{m})}_{\text{1st layer}}\underbrace{p(\boldsymbol{\gamma}_{m}|\mathbf{s}_{m})}_{\text{2nd layer}}\underbrace{p(\mathbf{s}_{m})}_{\text{3rd layer}},$ (25)
where the first layer is assigned a conditional Gaussian prior $p(\mathbf{x}_{m}|\boldsymbol{\gamma}_{m})=\mathcal{CN}\left(\mathbf{x}_{m};\mathbf{0},\mathbf{R}_{m}\right)$ with $\mathbf{R}_{m}=\mathrm{diag}\left(\boldsymbol{\gamma}_{m}^{-1}\right)$. The second layer $p(\boldsymbol{\gamma}_{m}|\mathbf{s}_{m})$ is modeled as a conditional Gamma distribution given by
$p(\boldsymbol{\gamma}_{m}|\mathbf{s}_{m})$ $=\prod_{n=1}^{N}p(\gamma_{m,n}|{s}_{m,n}),$ (26)
@@
$+\delta(\gamma_{m,n})\delta(1+s_{m,n}),$
where $s_{m,n}\in\left\{-1,1\right\}$ is a hidden state variable of variance precision parameter $\gamma_{m,n}$; $\delta(\cdot)$ denotes the Dirac Delta function; $a_{m}$ and $b_{m}$ denote the shape parameters of Gamma distribution. The third layer is assigned a MRF given by
$p(\mathbf{s}_{m})=\left(\prod_{n=1}^{N}\prod_{k\neq n}^{N}\psi(s_{m,n},s_{m,k})\right)^{\frac{1}{2}}\prod_{n=1}^{N}\phi(s_{m,n}),$ (28)
where $\psi(s_{m,n},s_{m,k})=\exp(\beta s_{m,n}s_{m,k})$ and $\psi(s_{m,n})=\exp(-\alpha s_{m,n})$ denote the pair-potential and self-potential functions with $\alpha$ and $\beta$ as hyperparameters.
@@
The joint posterior probability distribution $p(\mathbf{x}_{m}^{(t)},\boldsymbol{\gamma}_{m}^{(t)},\mathbf{s}_{m}^{(t)}|\boldsymbol{\mu}_{m}^{(t)})$ can be factorized as a pair-wise MRF, which is given by (30), as shown in the top of the next page, where $\mathbf{v}_{m,n}^{(t)}$ and $v^{n,k}_{m}(t)$ indicate the $n$-th column and $(n,k)$-th entry of $\mathbf{V}_{m}^{(t)}=(\boldsymbol{\Sigma}_{m}^{(t)})^{-1}$, respectively; $\psi_{\mathrm{post}}(x_{m,n}^{(t)})$ and $\phi_{\mathrm{post}}(x_{m,n}^{(t)},x_{m,k}^{(t)})$ denote the self potential of $x_{m,n}^{(t)}$, $\forall n\in\mathcal{N}$ and pair-potential functions between $x_{m,n}^{(t)}$ and $x_{m,k}^{(t)}$, $\forall n,k\in\mathcal{N}$ and $k\neq n$, respectively.
$p\left(\mathbf{x}_{m}^{(t)},\boldsymbol{\gamma}_{m}^{(t)},\mathbf{s}_{m}^{(t)}|\boldsymbol{\mu}_{m}^{(t)}\right)$ $\propto\prod_{n=1}^{N}\psi_{\mathrm{post}}\left(x_{m,n}^{(t)}\right)\prod_{n=1}^{N}\prod_{k\neq n}^{N}\phi_{\mathrm{post}}\left(x_{m,n}^{(t)},x_{m,k}^{(t)}\right),$ (30)
$\psi_{\mathrm{post}}\left(x_{m,n}^{(t)}\right)$ $\triangleq\psi_{\mathrm{like}}\left((\boldsymbol{\lambda}_{m}^{(t)})^{\mathrm{H}}\mathbf{v}_{m,n}^{(t)},v^{n,n}_{m}(t),x_{m,n}^{(t)}\right)\psi_{\mathrm{pri}}\left(\alpha_{m},x_{m,n}^{(t)}\right),$
$\phi_{\mathrm{post}}\left(x_{m,n}^{(t)},x_{m,k}^{(t)}\right)$ $\triangleq\phi_{\mathrm{like}}\left(v^{n,k}_{m}(t),x_{m,n}^{(t)},x_{m,k}^{(t)}\right)\phi_{\mathrm{pri}}\left(\beta_{m},x_{m,n}^{(t)},x_{m,k}^{(t)}\right).$

### v4

@@ II-B Problem Description
where $\varrho>0$ denotes a regularization factor, $\mathbf{H}_{\mathrm{T}}$ represents the angular-delay channel, and $\mathbf{F}_{\mathrm{A}}\in\mathbb{C}^{N_{{t}}\times N_{{t}}}$ and $\mathbf{F}_{\mathrm{D}}\in\mathbb{C}^{K\times K}$ are the discrete Fourier transform (DFT) matrices of dimensions $N_{{t}}$ and $K$, respectively. Up to now, several algorithms have been developed to address the problem in (9), including the SOMP method [32] and Bayesian inference-based approaches [33, 34, 35]. However, with the increasing number of antennas and subcarriers in mmWave XL-MIMO systems, these centralized algorithms become computationally prohibitive [24]. For example, the SBL algorithm incurs a computational complexity of $\mathcal{O}(M^{3}Q^{3}K^{3})$ due to the need for matrix inversions, making it unsuitable for large-scale scenarios.
@@ III Proposed Two-Stage Estimation Scheme
In this section, we propose a two-stage reconstruction-then-refinement estimation scheme, as illustrated in Fig. 2.
@@ III-A Stage I: Local Sparse Channel Reconstruction
$\mathbf{H}_{m}=\mathbf{F}_{\mathrm{L}}^{\mathrm{H}}\mathbf{H}_{\mathrm{T},m}^{\mathrm{L}}\mathbf{F}_{\mathrm{D}},m\in\mathcal{M},$ (10)
@@
$\mathbf{y}_{m}=\boldsymbol{\Phi}_{m}\mathbf{x}_{m}+\mathbf{n}_{m},m\in\mathcal{M},$ (11)
where $\mathbf{y}_{m}=\mathrm{vec}({\mathbf{Y}_{m}})\in\mathbb{C}^{QK\times 1}$ and $\mathbf{x}_{m}=\mathrm{vec}\left(\mathbf{H}_{\mathrm{T},m}^{\mathrm{L}}\right)\in\mathbb{C}^{NK\times 1}$ are the stacked received signal and local angular-delay channel, respectively. ${\boldsymbol{\Phi}}_{m}=\mathbf{F}_{\mathrm{D}}^{\mathrm{T}}\otimes\left(\mathbf{W}_{m}\mathbf{F}_{\mathrm{L}}^{\mathrm{H}}\right)\in\mathbb{C}^{QK\times NK}$ denotes the equivalent measurement matrix. Additionally, $\mathbf{n}_{m}=\mathrm{vec}({\mathbf{N}_{m}})\sim\mathcal{CN}(\mathbf{n}_{m};0,\zeta_{m}^{-1}\mathbf{I}_{QK})$ denotes the noise vector with covariance $\zeta_{m}^{-1}\mathbf{I}_{QK}$, where $\zeta_{m}$ is the unknown noise precision parameter to be estimated as well. Without loss of generality, a conjugated prior probability distribution is assigned for ${\zeta}_{m}$, i.e., $p(\zeta_{m})\propto\zeta_{m}^{-1}$.
@@
$\hat{\mathbf{x}}_{m}\propto\arg\max_{\mathbf{x}_{m}}\quad p(\mathbf{y}_{m}|\mathbf{x}_{m},\zeta_{m})p(\mathbf{x}_{m})p(\zeta_{m}),m\in\mathcal{M},$ (12)
where $p(\mathbf{y}_{m}|\mathbf{x}_{m},\zeta_{m})$ and ${p(\mathbf{x}_{m})}$ denote the joint likelihood function of $\mathbf{x}_{m}$ and $\zeta_{m}$ and the prior probability distribution of $\mathbf{x}_{m}$, respectively. Notably, the estimation performance of problem (12) is dependent on the accuracy of prior distribution $p(\mathbf{x}_{m})$. Meanwhile, due to the limited computational capability of LPUs, a computationally efficient estimation scheme is needed. To solve this problem, we proposed a SBL-GNNs algorithm, which will be elaborated on in the Sec. IV.
@@ III-B1 Channel Fusion in Angular Domain
The purpose of the central fusion is to align the local estimate $\hat{\mathbf{H}}_{\mathrm{T},m}^{\mathrm{L}}$ into the global angular domain with enhanced spatial resolution, thereby facilitating to leverage the correlations of subarray channel in the angular-delay domain. Specifically, leveraging the linear property of the DFT transformation matrix, we have
$\mathbf{H}_{\mathrm{T}}=\sum_{m=1}^{M}\mathbf{F}_{\mathrm{A},m}\mathbf{H}_{m}\mathbf{F}_{\mathrm{D}}^{\mathrm{H}}\triangleq\sum_{m=1}^{M}\mathbf{H}_{\mathrm{T},m},m\in\mathcal{M},$ (13)
where $\mathbf{H}_{\mathrm{T},m}$ denotes the angular-delay channels of subarray $m$ under the global angular transformation matrix $\mathbf{F}_{\mathrm{A}}=\left[\mathbf{F}_{\mathrm{A},1},\mathbf{F}_{\mathrm{A},2},\cdots,\mathbf{F}_{\mathrm{A},M}\right]$. Based on this, the aggregated channel in the CPU can be given by
@@
Compared to the local angular-delay channels $\mathbf{H}_{\mathrm{T},m}^{\mathrm{L}}$, the global angular-delay channel $\mathbf{H}_{\mathrm{T}}$ can more efficiently capture the inherent correlations among subarrays. Specifically, as the subarrays are spatially close to one another, $\mathbf{H}_{\mathrm{T},m}$ shares common angular and delay sparse supports, and exhibits high correlation. By effectively exploiting these common sparse supports and correlations, the reconstruction performance can be significantly enhanced.
@@ III-B2 Bayesian Refinement
where $p(\mathbf{r}|\mathbf{x},\kappa)$ denotes the joint likelihood function of $\mathbf{x}$ and $\kappa$; $p(\mathbf{x})$ and $p(\kappa)\propto\kappa^{-1}$ denote the prior probability distribution of $\mathbf{x}$ and $\kappa$, respectively. To solve problem (16) efficiently, we propose a variational message-passing-based denoising algorithm that leverages a Markov chain-based hierarchical prior to capture the sparsity and correlation of $\mathbf{x}$, as detailed in the Sec. V. Once the global angular-delay channel $\tilde{\mathbf{x}}$ is obtained, the global antenna-frequency-domain channel can be reconstructed as $\tilde{\mathbf{H}}=\mathbf{F}_{\mathrm{A}}^{\mathrm{F}}\tilde{\mathbf{H}}_{\mathrm{T}}\mathbf{F}_{\mathrm{D}}$ = $[\tilde{\mathbf{H}}_{1}^{\mathrm{T}},\tilde{\mathbf{H}}_{2}^{\mathrm{T}},\cdots,\tilde{\mathbf{H}}_{M}^{\mathrm{T}}]$. Then, $\tilde{\mathbf{H}}_{m}$ are sent to the corresponding LPUs for subsequent algorithm designs.
Remark 2.
Owing to its reconstruction-then-refinement strategy, the proposed two-stage estimation scheme achieves accuracy comparable to centralized methods while maintaining low computational complexity. Specifically, in the first stage, high-dimensional matrix inversion and matrix-vector multiplications are decomposed into several parallel low-dimensional operations, as detailed in Sec. IV. In the second stage, only scalar operations are involved, as discussed in Sec. V. Consequently, the proposed approach significantly reduces computational complexity compared to traditional centralized estimation schemes. Moreover, the centralized refinement step effectively exploits correlations across subchannels from different LPUs, further enhancing estimation performance and bringing it closer to that of centralized estimation.
@@ IV-A SBL Estimator
$p(\mathbf{x}_{m}|\boldsymbol{\gamma}_{m})$ $=\prod_{n\in\mathcal{V}}\mathcal{CN}(x_{m,n};0,\gamma_{m,n}^{-1}),m\in\mathcal{M}$ (17)
$p(\boldsymbol{\gamma}_{m})$ $=\prod_{n\in\mathcal{V}}p(\gamma_{m,n};\boldsymbol{\lambda}_{m,n}),m\in\mathcal{M},$ (18)
where $x_{m,n}$ and $\gamma_{m,n}$ are the $n$-th elements of $\mathbf{x}_{m}$ and $\boldsymbol{\gamma}_{m}$ corresponding to subarray $m$ with $n\in\mathcal{V}$, and $\gamma_{m,n}^{-1}$ denote the prior variance. In addition, $\boldsymbol{\lambda}_{m,n}$ denote the prior distribution parameters of ${\gamma}_{m,n}$. Utilizing (17), (18) and Expectation-Maximization (EM) framework, problem (12) can be efficiently solved by SBL-based algorithms [34, 35], and the update steps in each iteration are as follows:
@@
$\boldsymbol{\lambda}_{m,n}^{(t)}$ $=g_{2}\left(\boldsymbol{\mu}_{m}^{(t-1)},\boldsymbol{\Sigma}_{m}^{(t-1)},\boldsymbol{\gamma}_{m}^{(t)}\right),$ (22)
@@
where $\boldsymbol{\Sigma}_{m}^{(t)}$ and $\boldsymbol{\mu}_{m}^{(t)}$ represent the posterior covariance matrix and mean of $\mathbf{x}_{m}$ at the $t$-th iteration, respectively; $g_{1}(\cdot)$, $g_{2}(\cdot)$, and $g_{3}(\cdot)$ denote the estimators for the prior variance, the hyperparameters of the prior variance, and the noise precision parameters, respectively. Specifically, $g_{1}(\cdot)$ leverages the $(t-1)$-th estimation results from the observation module and the $(t-1)$-th hyperparameters of the prior variance to update the $t$-th variance estimation. Subsequently, using the $(t-1)$-th estimation results from the observation module along with the current variance estimation, the hyperparameters of the prior variance are updated. Finally, the noise precision parameters are refined based on the inputs from the observation module and the observed signal. These updates are performed iteratively until convergence, at which point the posterior mean $\boldsymbol{\mu}_{m}^{(t)}$ is adopted as the final estimate of $\mathbf{x}_{m}$.
Recently, various approaches have been proposed for designing $g_{1}(\cdot)$, $g_{2}(\cdot)$, and $g_{3}(\cdot)$, including mean-field-based methods [34] and variational inference techniques [35]. However, conventional sparse prior models often fail to fully leverage the correlations among channel entries. For instance, the Bernoulli-Gaussian prior [36] and two-layer nested Gaussian-Gamma prior [37] assume independence among channel entries, which leads to performance degradation when addressing block-sparse signals. Although advanced approaches such as the Markov chain-based and the 4-connected MRF-based hierarchical priors [14, 15, 34, 35] have been developed, these methods only consider correlations among adjacent entries. This limitation renders them inadequate for effectively capturing the global correlations of local angular-delay channels.
@@ IV-B Proposed SBL-GNNs Estimator
To overcome the limitations of conventional sparse prior models, we propose a GNN-based approach within the SBL framework to estimate the prior variance of the angular-delay channel and its associated parameters. Specifically, the sparse prior of $\mathbf{x}_{m}$ is modeled as a fully connected MRF. Leveraging the structural properties of the pairwise MRF, a GNN architecture is designed to replace the non-linear mapping functions $g_{1}(\cdot)$ and $g_{2}(\cdot)$ in the M-step are replaced with GNNs, thereby efficiently capturing the dependencies among channel coefficients. Consequently, the proposed SBL-GNNs algorithm can be regarded as a deep-unfolding counterpart of the traditional SBL algorithm with a fully connected MRF-based sparse prior.
According to SBL principles, the output of the observation module in the $t$-th iteration can be equivalently characterized as an AWGN model, i.e.,
$\boldsymbol{\mu}_{m}^{(t)}=\mathbf{x}_{m}+\mathbf{w}_{m}^{(t)},m\in\mathcal{M},$ (24)
@@
$p(\mathbf{x}_{m},\boldsymbol{\gamma}_{m};\mathbf{s}_{m})$ $={p(\mathbf{x}_{m}|\boldsymbol{\gamma}_{m})}{p(\boldsymbol{\gamma}_{m}|\mathbf{s}_{m})}{p(\mathbf{s}_{m})},$ (25)
where the first layer is assigned a conditional Gaussian prior $p(\mathbf{x}_{m}|\boldsymbol{\gamma}_{m})=\mathcal{CN}\left(\mathbf{x}_{m};\mathbf{0},\mathbf{R}_{m}\right)$ with $\mathbf{R}_{m}=\mathrm{diag}\left(\boldsymbol{\gamma}_{m}^{-1}\right)\in\mathbb{R}^{NK\times NK}$. The second layer $p(\boldsymbol{\gamma}_{m}|\mathbf{s}_{m})$ is modeled as a conditional Gamma distribution given by
$p(\boldsymbol{\gamma}_{m}|\mathbf{s}_{m})$ $=\prod_{n\in\mathcal{V}}p(\gamma_{m,n}|{s}_{m,n}),m\in\mathcal{M},$ (26)
@@
$+\delta(\gamma_{m,n})\delta(1+s_{m,n}),m\in\mathcal{M},n\in\mathcal{V},$
where $s_{m,n}\in\left\{-1,1\right\}$ denotes the $n$-th element of $\mathbf{s}_{m}$ is a hidden state variable of variance precision parameter $\gamma_{m,n}$ with $m\in\mathcal{M}$ and $n\in\mathcal{V}$; $\delta(\cdot)$ denotes the Delta function; $a_{m}$ and $b_{m}$ denote the shape parameters of Gamma distribution. Finally, the third layer is assigned a MRF given by
$p(\mathbf{s}_{m})=\left(\prod_{n\in\mathcal{V}}\prod_{k\neq n}\psi(s_{m,n},s_{m,k})\right)^{\frac{1}{2}}\prod_{n\in\mathcal{V}}\phi(s_{m,n}),$ (28)
where $\psi(s_{m,n},s_{m,k})=\exp(\beta_{m}s_{m,n}s_{m,k})$ and $\psi(s_{m,n})=\exp(-\alpha_{m}s_{m,n})$, $m\in\mathcal{M}$, $n,k\in\mathcal{V}$ denote the pair-potential and self-potential functions with $\alpha_{m}$ and $\beta_{m}$ as hyperparameters of MRF.
@@
The joint posterior probability distribution $p(\mathbf{x}_{m},\boldsymbol{\gamma}_{m},\mathbf{s}_{m}|\boldsymbol{\mu}_{m}^{(t)})$ can be factorized as a pair-wise MRF, which is given by (29), as shown in the top of this page, where $\mathbf{v}_{m,n}^{(t)}$ and $v^{n,k}_{m}(t)$ indicate the $n$-th column and $(n,k)$-th entry of $\mathbf{V}_{m}^{(t)}=(\boldsymbol{\Sigma}_{m}^{(t)})^{-1}$, respectively; $\psi_{\mathrm{post}}(x_{m,n})$ and $\phi_{\mathrm{post}}(x_{m,n},x_{m,k})$ denote the self potential of $x_{m,n}$, $\forall n\in\mathcal{V}$ and pair-potential functions between $x_{m,n}$ and $x_{m,k}$, $\forall n,k\in\mathcal{V}$ and $k\neq n$, respectively.
$p\left(\mathbf{x}_{m},\boldsymbol{\gamma}_{m},\mathbf{s}_{m}|\boldsymbol{\mu}_{m}^{(t)}\right)$ $\propto\prod_{n\in\mathcal{V}}\psi_{\mathrm{post}}\left(x_{m,n}\right)\prod_{n\in\mathcal{V}}\prod_{k\neq n}\phi_{\mathrm{post}}\left(x_{m,n},x_{m,k}\right),m\in\mathcal{M},n,k\in\mathcal{V},$ (29)
$\psi_{\mathrm{post}}\left(x_{m,n}\right)$ $\triangleq\psi_{\mathrm{like}}\left((\boldsymbol{\lambda}_{m}^{(t)})^{\mathrm{H}}\mathbf{v}_{m,n}^{(t)},v^{n,n}_{m}(t),x_{m,n}\right)\psi_{\mathrm{pri}}\left(\alpha_{m},x_{m,n}\right),$
$\phi_{\mathrm{post}}\left(x_{m,n},x_{m,k}\right)$ $\triangleq\phi_{\mathrm{like}}\left(v^{n,k}_{m}(t),x_{m,n},x_{m,k}\right)\phi_{\mathrm{pri}}\left(\beta_{m},x_{m,n},x_{m,k}\right).$